pls. help me solve this problem... thank you
a normal distribution has the mean of 160.5. if 75% of the area under the curve lies to the left of 185.2 find:
a. the area to the left of 160.5
b. the area to the right of 185.2
c. the area between 160.5 and 185.2
d. the standard deviation
e. the area betweem 160.5 and 180.5
f. the area to the right of 180.5

Answer by ewatrrr(24785)


Hi
a normal distribution has the mean of 160.5.
if 75% of the area under the curve lies to the left of 185.2
a. the area to the left of 160.5(μ) is 50%
b. the area to the right of 185.2 is 25%
c. the area between 160.5 and 185.2 is 25%
d. the standard deviation: z(P =.75) = .67449 = (185.2-160.5)/σ σ = 36.62
e. the area betweem 160.5 and 180.5: z = 20/36.62 = .5461 and P =.7075
A = 70.75% - 50% = 20.75%
f. the area to the right of 180.5 is 29.25%
Below... left of the blue line, is 75% of the area under the curve